# A self-supervised framework for learning whole slide representations

### v1

@@ 5.4 Self-attention visualizations reveal tissue phenotypes in full gigapixel WSIs
We evaluate S3L qualitatively by visualizing the self-attention on the full WSI. As illustrated in Figure 6, self-attention maps on H&E WSI can distinguish different tissue phenotypes such as blood, dense tumor, and necrosis. The CLS token attends to dense tumor regions instead of non-diagnostic regions such as blood and necrosis for slide representations. This observation is consistent with existing literature on self-supervised vision transformers, where attention maps can serve as unsupervised segmentation (Caron et al., 2021; Chen et al., 2022b). To our knowledge, our work is the first to generate self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions, showing the strong representation capacity, flexibility, and efficiency of S3L. More H&E and SRH attention map visualizations are in Appendix C.4.
@@ 6 Conclusion
We present S3L, a transformer-based SSL approach for WSI representation learning, in a general, versatile, and lightweight framework. S3L uses domain-informed, vision-language transformations for high-quality WSI view generation and self-supervised learning. We demonstrated that S3L achieved superior performance on two computational pathology benchmarks and datasets, cancer diagnosis and molecular genetic prediction. To our knowledge, this is the first systematic study of whole slide representation learning evaluating the role of SSL objectives, transformation strategies, and patch encoders.
Limitations
Our experimentation is limited to two biomedical microscopy modalities and three major self-supervised learning paradigms. Future work includes validating S3L on additional datasets, including H&E from different organ systems, or fluorescent microscopy, as well as different SSL paradigms such as generative SSL. Another limitation of S3L is hyperparameter tuning for different configurations of SCM transformation, which are domain and dataset-specific. The future direction of our work also includes efficient end-to-end training to unify the training of patch and slide encoders.
Impact Statement
This paper presents with the aim to advance the field of whole slide representation learning and computational pathology. Our work builds toward an automated diagnostic support system for biomedical microscopy and histopathology, as it paves the way for more accurate diagnoses and personalized treatment recommendations. Moreover, we strive to develop self-supervised learning frameworks in order to promote and maintain data privacy and ethical data use of patient data within healthcare AI systems. Additionally, SSL is essential for developing medical foundation models. We hope that S3L can serve as a potential foundation model training strategy within computational pathology. While our experiments are in medical applications, it has a potential impact on other similar gigapixel image modalities, such as geospatial and astronomical images.
Acknowledgement
We would like to thank Karen Eddy and Lin Wang for their administrative support and data collection.
This work is supported in part by grants NIH T32GM141746, NIH F31NS135973, NIH K12NS080223, NIH R01CA226527, Cook Family Brain Tumor Research Fund, Mark Trauner Brain Research Fund: Zenkel Family Foundation, Ian’s Friends Foundation, and the Investigators Awards grant program of Precision Health at the University of Michigan.
@@ Appendix A Data Details
SRH dataset.
Our SRH benchmark includes common brain tumors and normal brain tissue histology. Our data collection and processing follows the protocol described in (Jiang et al., 2022), where each specimen is scanned with a commercially available stimulated Raman histology microscope manufactured by Invenio Imaging. Each specimen is diagnosed by a board-certified neuropathologist and a detailed number of slides per class for SRH is in Table 4.
@@
Table 4: SRH dataset breakdown. Number of slides in the training and validation set for supervised training and KNN evaluation. HGG, high grade glioma, LGG, low grade glioma, mening., meningioma, metast., metastasis, pit., pituitary adenoma, schwan., schwannoma, normal, normal brain tissue.
H&E dataset.
Our H&E benchmark includes glioma specimens from the Cancer Genome Atlas Program (TCGA) and the Digital Brain Tumour Atlas (DBTA) (Roetzer-Pejrimovsky et al., 2022). These publicly available datasets feature H&E images collected in both the United States and Europe. Slides are divided into three classes based on molecular labels as defined by the World Health Organization (Louis et al., 2021), using IDH and 1p/19q co-deletion status included in both datasets. At training time, both DBTA and TCGA training set is used, and a separate, held-out validation set from the TCGA dataset is used for benchmarking. A detailed number of slides per class is in Table 5. Our H&E data processing pipeline follows (Jiang et al., 2023). Each WSI is divided into 300x300 patches, and blank / background regions are excluded. All patches used in training are stain-normalized using the Macenko algorithm (Macenko et al., 2009).
@@
Oligo. IDH-mut, 1p/19q-codel 176 265 62
Astro. IDH-mut 157 360 88
GBM IDH-wt 619 732 191
Table 5: H&E dataset breakdown. Number of slides in the training and validation set from both TCGA and DBTA. Oligo, oligodendroglioma, Astro, astrocytoma, GBM, glioblastoma.
@@ Appendix B Methods and Implementation
B.1 SCM transformation details
As described in section 3.3, we implement our SCM transformations with pre-computed embeddings and patch coordinates. Recall that each whole slide is represented as a $(\boldsymbol{x},\boldsymbol{p})$ tuple, where $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ is the embedding tensor and $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$ is a corresponding coordinate tensor for each patch in the WSI. Each transformation is implemented by indexing into the rows (corresponding to patches/tokens) of $\boldsymbol{x}$ and $\boldsymbol{p}$. A PyTorch style pseudocode is in Algorithm 1.
@@ B.1 SCM transformation details
Algorithm 1 S3L transformations in PyTorch style
@@ B.2 Relative positional embedding
To handle the different sizes of WSIs, it is non-trivial to use a fixed-size learnable position embedding with absolute distance (Devlin et al., 2018). We adopt the relative positional encoding introduced in (Li et al., 2021b), which utilizes the learnable Fourier feature, modulated with a multi-layer perceptron. This positional encoding takes the coordinates of the patch encoder as input and the output is added directly to the transformer with the same dimension. We visualize the similarity of positional embeddings in Figure 8.
Figure 8: Cosine similarity of learned positional embeddings. Visualization is generated similarly as in (Dosovitskiy et al., 2020): each patch shows the cosine similarity between the position embedding of the token with the indicated row and column and the position embeddings of all other tokens.
@@ Appendix C Extended Results
C.1 S3L results with ImageNet and HiDisc patch encoders
In Figure 5, we compared the S3L’s performance to the mean pooling baseline and supervised SCM performance upper bound. We present the detailed benchmark results for ImageNet and HiDisc patch encoders in Tables 6 and 7, respectively. S3L significantly improves slide representations in all self-supervised learning paradigms experimented, for both ImageNet and HiDisc patch encoders. Notably, in our H&E benchmark, S3L slide representations trained with the ImageNet patch encoder outperform pooling results with the HiDisc patch encoder, demonstrating S3L’s versatility in learning slide representations with out-of-distribution patch encoders.
SRH H&E
MCA F1 MCA F1
Max pooling 53.5 53.7 57.4 57.6
Mean pooling 53.1 53.0 59.8 74.2
S3L-SimCLR 71.7 (0.7) 73.2 (0.4) 74.0 (1.9) 73.0 (1.9)
S3L-VICReg 66.2 (0.7) 66.6 (0.7) 71.7 (0.7) 71.0 (1.1)
S3L-BYOL 69.3 (1.6) 69.9 (1.2) 75.3 (2.5) 74.0 (2.5)
Supervised SCM 75.4 (1.4) 76.1 (0.8) 76.4 (1.1) 76.4 (1.2)
Table 6: S3L results with ImageNet patch encoder. Each experiment is run 3 times, and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
SRH H&E
MCA F1 MCA F1
Max pooling 82.3 83.5 73.3 73.4
Mean pooling 82.7 82.9 74.9 60.1
S3L-SimCLR 83.4 (0.1) 82.9 (0.3) 77.5 (1.3) 77.1 (1.3)
S3L-VICReg 84.0 (0.7) 83.9 (0.7) 75.3 (1.7) 75.0 (1.7)
S3L-BYOL 84.2 (1.4) 83.9 (1.2) 75.4 (2.1) 75.0 (2.0)
Supervised SCM 87.4 (0.8) 87.5 (0.7) 83.6 (0.3) 82.4 (0.2)
Table 7: S3L results with HiDisc patch encoder. Each experiment is run 3 times and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
Patch encoder Slide encoder SRH H&E
MCA F1 MCA F1
HIPT Max pooling 47.5 47.2 58.5 59.0
Mean pooling 55.1 56.4 54.7 55.0
S3L-SimCLR 70.1 (0.4) 71.9 (0.5) 74.9 (1.7) 73.9 (1.7)
Supervised SCM 72.5 (1.1) 73.7 (0.9) 76.6 (1.8) 76.1 (1.6)
PLIP Max pooling 50.2 50.4 63.3 62.6
Mean pooling 51.7 51.9 66.8 66.3
S3L-SimCLR 68.2 (0.4) 69.7 (0.4) 76.2 (2.0) 75.5 (2.2)
Supervised SCM 72.1 (1.4) 72.6 (1.2) 82.6 (0.7) 81.5 (0.9)
Table 8: Out of distribution S3L benchmarks. Each experiment is run 3 times and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
C.2 S3L results with out-of-distribution patch encoders
In Figure 5, we also benchmarked the S3L’s performance when trained with HIPT (Chen et al., 2022b) and PLIP (Huang et al., 2023) patch encoders. We present the detailed benchmarks for these experiments with F1 score in Table 8. Recall that the HIPT patch encoder is trained with the complete TCGA dataset, and the PLIP patch encoder is trained with publicly available (non-WSI) histology images on Twitter. For SRH, HIPT and PLIP patch encoders are OOD, and for H&E, HIPT and PLIP are near domain.
C.3 Extended ablation results

### v2

@@ 5.5 Self-attention visualizations reveal tissue phenotypes in full gigapixel WSIs
Finally, we adopted the strategy in [47] and evaluated ssSPT qualitatively by visualizing self-attention on the full WSI. As illustrated in Figure 6, self-attention maps on H&E WSI can distinguish different tissue phenotypes such as blood, dense tumor, and necrosis. The CLS token attends to dense tumor regions instead of non-diagnostic regions such as blood and necrosis for slide representations. This observation is consistent with existing literature on self-supervised vision transformers, where attention maps can serve as unsupervised segmentation [47, 5]. To our knowledge, our work is the first to generate self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions, showing the strong representation capacity, flexibility, and efficiency of ssSPT. More H&E and SRH attention map visualizations are in Appendix C.6.
Figure 6: Attention heatmap visualization. Self-attention visualization demonstrates ssSPT’s ability to distinguish between different tissue phenotypes such as blood, dense tumor, and necrosis. To our knowledge, these are the first self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions.
@@ 6 Conclusion
We present Slide Pre-trained Transformers (SPT), a self-supervised approach for whole slide representation learning, in a general, versatile, and lightweight framework. SPT uses domain-informed, vision-language transformations for high-quality WSI view generation and self-supervised learning. We demonstrated that SPT achieves superior performance on five computational pathology benchmarks across three clinical tasks, including histopathologic diagnosis, cancer subtyping, and molecular genetic prediction. To our knowledge, this is the first systematic study of whole slide representation learning evaluating the role of SSL objectives, transformation strategies, and patch encoders.
Limitations.
Our experimentation was limited to two biomedical microscopy modalities and three major self-supervised learning paradigms. Future work includes validating SPT on additional datasets, including WSIs from different microscopy modalities, such as fluorescent microscopy, as well as different SSL paradigms such as generative SSL. Another limitation of SPT is hyperparameter tuning for different configurations of SPT transformations, which are domain and dataset-specific. The future direction of our work also includes efficient end-to-end training to unify the training of patch and slide encoders.
Broader impact.
This paper aims to advance the field of whole slide representation learning and computational pathology. Our work builds toward an automated diagnostic support system for biomedical microscopy and histopathology, paving the way for more accurate diagnoses and personalized treatment recommendations. Moreover, we strive to develop self-supervised learning frameworks in order to promote and maintain data privacy and ethical use of patient data within healthcare AI systems. Additionally, SSL is essential for developing medical foundation models. We hope that SPT can serve as a potential foundation model training strategy within computational pathology. While our experiments are in medical applications, there is a potential for impact on other similar gigapixel image modalities, such as geospatial and astronomical images.
Acknowledgements and Competing Interests
We would like to thank Karen Eddy and Lin Wang for their administrative support and data collection efforts.
This work was supported, in part, by the National Institutes of Health (NIH) grants F31NS135973 (C.J.), T32GM141746 (C.J.), and K12NS080223 (T.H.). This work was also supported, in part, by the Chan Zuckerberg Foundation (CZI) Advancing Imaging Through Collaborative Project grant (T.H.), the Cook Family Brain Tumor Research Fund (T.H.), the Mark Trauner Brain Research Fund (T.H.), the Zenkel Family Foundation (T.H.), Ian’s Friends Foundation (T.H.) and the UM Precision Health Investigators Awards grant program (T.H.).
@@ Appendix A Data Details
The five benchmarks used in our experimentation are described in Table 4.
Benchmark # clases Tissue organ Imaging modality Task
SRH CNS 7 Brain SRH Histological classification
H&E Glioma 3 Brain H&E stained Molecular classification
TCGA BRCA 2 Breast H&E stained Tumor subtyping
TCGA NSCLC 2 Lung H&E stained Tumor subtyping
TCGA RCC 3 Kidney H&E stained Tumor subtyping
Table 4: Benchmark overview. Our benchmarks span four different organs and two different microscopy modalities.
A.1 SRH CNS benchmark
Our SRH benchmark includes common brain tumors and normal brain tissues. Our data collection and processing follows the protocol in [51], where each specimen is scanned with a commercially available stimulated Raman histology microscope manufactured by Invenio Imaging. Each specimen is diagnosed by a board-certified neuropathologist. The number of slides in each class is in Table 5.
@@ A.1 SRH CNS benchmark
Table 5: SRH dataset breakdown. Number of slides in the training and validation set for supervised training and evaluation. HGG, high grade glioma, LGG, low grade glioma, mening., meningioma, metast., metastasis, pit., pituitary adenoma, schwan., schwannoma, normal, normal brain tissue.
A.2 H&E Glioma benchmark
Our H&E benchmark includes glioma specimens from the Cancer Genome Atlas Program (TCGA) and the Digital Brain Tumour Atlas (DBTA) [52]. These publicly available datasets feature H&E images collected in both the United States and Europe. Slides are divided into three classes based on molecular labels as defined by the World Health Organization [53], using IDH and 1p/19q co-deletion status included in both datasets. At training time, both DBTA and TCGA training set is used, and a separate, held-out validation set from the TCGA dataset is used for benchmarking. A detailed number of slides per class is in Table 6. Our H&E data processing pipeline follows [13]. Each WSI is divided into 300 $\times$ 300 patches, and blank / background regions are excluded. All patches used in training are stain-normalized using the Macenko algorithm [56].
@@ A.2 H&E Glioma benchmark
Oligodendroglioma IDH mutant, 1p/19q co-deleted 176 265 62
Astrocytoma IDH mutant 157 360 88
Glioblastoma IDH wildtype 619 732 191
Table 6: H&E glioma dataset breakdown. Number of slides from both datasets.
@@ Appendix B Methods and Implementation
B.1 SPT transformation details
As described in section 3.3, we implement our SPT transformations with pre-computed embeddings and patch coordinates. Recall that each whole slide is represented as a $(\boldsymbol{x},\boldsymbol{p})$ tuple, where $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ is the embedding tensor and $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$ is a corresponding coordinate tensor for each patch in the WSI. Each transformation is implemented by indexing into the rows (corresponding to patches/tokens) of $\boldsymbol{x}$ and $\boldsymbol{p}$. A PyTorch style pseudocode is in Algorithm 1.
@@ B.1 SPT transformation details
Algorithm 1 SPT transformations in PyTorch style
@@ B.2 Relative positional embedding
To work with WSIs of different sizes, it is non-trivial to use a fixed-size learnable position embedding with absolute distance [34]. We adopt the relative positional encoding introduced in [55], which utilizes the learnable Fourier feature, modulated with a multi-layer perceptron. This positional encoding uses the coordinates of the patch encoder as input and the output is added directly to the transformer with the same dimension. We visualize the positional embedding similarity in Figure 7.
Figure 7: Cosine similarity of learned positional embeddings. Visualization is generated similarly as in [57]: each patch shows the cosine similarity between the position embedding of the token with the indicated row and column and the position embeddings of all other tokens.
@@ Appendix C Extended Results
C.1 Extended ssSPT evaluation with kNN classifier
We evaluate ssSPT and baselines using a $k$NN classifier in Table 1. Error bars from the table are reported in Table 7.
Benchmark Method MCA F1 AUC
SRH CNS Mean pooling 72.5 73.2 94.8
Max pooling 73.5 75.6 95.0
Giga-SSL 71.2 (0.9) 72.7 (0.5) 94.2 (1.4)
ssSPT (Ours) 82.3 (0.5) 82.3 (0.3) 94.7 (0.1)
H&E Glioma Mean pooling 68.3 67.9 87.9
Max pooling 68.2 69.2 87.0
Giga-SSL 71.6 (1.3) 71.5 (1.1) 89.1 (0.2)
ssSPT (Ours) 76.5 (0.5) 76.1 (0.4) 90.9 (0.4)
TCGA BRCA Mean pooling 58.1 (4.0) 27.4 (11.3) 74.7 (7.9)
Max pooling 58.3 (3.8) 29.1 (10.1) 77.6 (10.1)
HIPT 63.2 (6.1) 39.9 (13.7) 77.5 (4.2)
Giga-SSL 70.4 (9.7) 53.9 (17.5) 85.4 (6.7)
ssSPT (Ours) 72.7 (4.9) 58.3 (6.8) 80.4 (7.8)
TCGA NSCLC Mean pooling 75.1 (3.7) 76.5 (3.4) 85.3 (2.9)
Max pooling 79.3 (3.9) 80.7 (3.2) 88.4 (1.7)
HIPT 80.6 (2.9) 80.1 (3.3) 88.9 (2.7)
Giga-SSL 84.3 (3.0) 84.9 (2.6) 92.3 (2.7)
ssSPT (Ours) 86.1 (2.3) 86.1 (2.6) 92.3 (1.7)
TCGA RCC Mean pooling 84.2 (3.3) 86.3 (2.8) 96.6 (1.0)
Max pooling 79.4 (4.5) 82.8 (3.7) 96.6 (1.7)
HIPT 88.4 (3.2) 89.2 (3.0) 97.4 (1.6)
Giga-SSL 90.0 (2.5) 89.3 (2.9) 97.5 (1.0)
ssSPT (Ours) 91.7 (2.1) 90.5 (2.7) 98.3 (1.3)
Table 7: Self-supervised benchmarks. We use kNN classifier to evaluate ssSPT and baselines. Extended Table 1 with standard deviations reported in (parentheses).
C.2 Extended ssSPT and suSPT evaluation with linear classifier
We evaluate ssSPT, suSPT, and baselines using linear evaluation in Table 2. KNN evaluation results and error bars from the table are reported in Table 8.
kNN Linear evaluation
Benchmark Method MCA F1 AUC MCA F1 AUC
SRH CNS Self-supervised methods
Giga-SSL 71.2 (0.9) 72.7 (0.5) 94.2 (1.4) 78.6 (1.8) 75.6 (2.1) 96.1 (0.3)
ssSPT (Ours) 82.3 (0.5) 82.3 (0.3) 94.7 (0.1) 85.4 (0.3) 83.2 (0.3) 97.6 (0.3)
Supervised methods
ABMIL 77.0 (0.9) 76.3 (0.1) 93.1 (0.6) 79.8 (0.4) 78.9 (1.0) 96.0 (0.4)
CLAM - - - 83.5 (1.3) 82.9 (1.2) 96.7 (0.3)
TransMIL 84.8 (0.8) 84.7 (0.3) 92.2 (0.1) 84.4 (1.2) 84.0 (0.7) 84.0 (0.7)
suSPT (Ours) 85.0 (0.5) 84.2 (0.6) 92.8 (0.7) 85.6 (0.5) 84.4 (0.5) 97.2 (0.3)
H&E Glioma Self-supervised methods
Giga-SSL 71.6 (1.3) 71.5 (1.1) 89.1 (0.2) 76.5 (0.6) 73.9 (0.4) 90.7 (0.3)
ssSPT (Ours) 76.5 (0.5) 76.1 (0.4) 90.9 (0.4) 80.0 (2.4) 78.0 (2.2) 91.9 (0.8)
Supervised methods
ABMIL 75.6 (1.8) 75.7 (1.8) 88.4 (0.4) 75.0 (0.6) 74.5 (0.6) 90.2 (0.3)
CLAM - - - 78.9 (2.2) 77.7 (1.8) 88.5 (1.5)
TransMIL 79.1 (2.5) 78.8 (2.7) 86.5 (0.3) 80.0 (1.4) 79.4 (1.8) 91.9 (1.1)
suSPT (Ours) 80.5 (1.6) 79.9 (1.4) 88.3 (0.5) 80.7 (2.4) 80.3 (1.9) 93.7 (0.3)
TCGA BRCA Self-supervised methods
Giga-SSL 70.4 (9.7) 53.9 (17.5) 85.4 (6.7) 81.7 (6.5) 62.6 (9.2) 91.2 (5.4)
ssSPT (Ours) 72.7 (4.9) 58.3 (6.8) 80.4 (7.8) 83.0 (7.3) 67.4 (9.9) 89.4 (6.3)
Supervised methods
ABMIL 71.7 (9.7) 56.0 (17.5) 84.9 (9.1) 78.6 (9.1) 62.8 (14.3) 86.1 (7.5)
CLAM - - - 76.8 (8.9) 59.3 (13.8) 85.8 (6.3)
DSMIL - - - 77.0 (6.5) 58.0 (7.8) 83.8 (7.0)
TransMIL 79.0 (8.1) 67.7 (11.2) 84.7 (6.8) 80.8 (8.1) 69.4 (11.3) 90.0 (5.5)
HIPT - - - 75.8 (8.2) 60.2 (11.7) 87.4 (5.7)
suSPT (Ours) 78.2 (9.4) 66.1 (17.1) 86.1 (6.1) 84.2 (5.3) 70.4 (8.2) 90.5 (4.8)
TCGA NSCLC Self-supervised methods
Giga-SSL 84.3 (3.0) 84.9 (2.6) 92.3 (2.7) 87.5 (2.3) 87.9 (2.2) 93.7 (2.2)
ssSPT (Ours) 86.1 (2.3) 86.1 (2.6) 92.3 (1.7) 88.0 (3.1) 88.0 (3.2) 94.8 (2.1)
Supervised methods
ABMIL 80.1 (4.0) 79.6 (4.5) 89.7 (2.1) 86.8 (2.4) 86.7 (2.3) 93.7 (1.2)
CLAM - - - 84.3 (3.4) 83.6 (4.1) 92.8 (2.0)
DSMIL - - - 84.9 (3.8) 84.0 (4.6) 92.0 (2.3)
TransMIL 88.8 (2.4) 88.9 (2.5) 94.2 (2.6) 87.4 (2.8) 87.5 (2.8) 95.3 (1.5)
HIPT - - - 88.2 (2.5) 88.2 (2.5) 95.2 (2.0)
suSPT (Ours) 90.3 (1.8) 90.4 (2.0) 94.1 (2.6) 90.3 (2.1) 90.4 (2.2) 95.7 (2.4)
TCGA RCC Self-supervised methods
Giga-SSL 90.0 (2.5) 89.3 (2.9) 97.5 (1.0) 89.9 (2.7) 87.8 (2.6) 97.7 (0.9)
ssSPT (Ours) 91.7 (2.1) 90.5 (2.7) 98.3 (1.3) 93.4 (1.7) 91.1 (2.1) 98.7 (1.1)
Supervised methods
ABMIL 89.5 (6.0) 90.6 (5.2) 97.5 (2.2) 89.9 (2.6) 89.4 (3.2) 97.8 (1.1)
CLAM - - - 85.6 (8.0) 83.6 (7.0) 97.3 (1.7)
DSMIL - - - 90.1 (3.7) 87.1 (4.4) 97.1 (1.5)
TransMIL 90.6 (2.3) 90.0 (2.5) 96.6 (1.3) 87.4 (2.8) 87.5 (2.8) 95.3 (1.5)
HIPT - - - 92.0 (2.0) 92.1 (2.6) 98.0 (1.2)
suSPT (Ours) 92.1 (3.0) 91.1 (3.5) 96.3 (2.1) 92.3 (2.9) 91.0 (3.5) 98.6 (0.7)
Table 8: SPT benchmarks. Extended Table 2 with KNN evaluation results and standard deviations. Standard deviations are reported in (parentheses).
C.3 ssSPT results with different patch encoders
In Figure 5(b), we compared the ssSPT’s performance to the pooling baseline and suSPT performance upper bound with the H&E glioma benchmark. Here, we present the same experiment with the SRH CNS benchmark in Figure 8. In the SRH benchmark, WSI representations improved dramatically in OOD and foundation patch encoders, which are not trained on SRH data. After training with SPT, as demonstrated by a maximum 18.6% boost in MCA for ImageNet. We also observe the expected performance boost for in-domain patch encoder SimCLR and HiDisc. Similar to the H&E glioma benchmark, ssSPT performance approaches suSPT upperbound on all patch encoders, within 4 points in MCA for SRH CNS benchmarks.
Figure 8: SPT results on SRH CNS with a wide range of patch encoders. Publicly available foundation patch encoders are trained using H&E data and are out-of-distribution (OOD) for SRH. In-domain patch encoders are trained using SRH data.
In addition, full metrics for each patch encoder for both SRH CNS and H&E glioma are in Table 9.
SRH H&E Gliomas
Patch Encoder Method MCA F1 AUC MCA F1 AUC
Pooling 53.1 53.0 88.1 59.8 74.2 81.7
ImageNet ssSPT 71.7 (0.7) 73.2 (0.4) 93.7 (0.3) 75.3 (2.5) 74.0 (2.5) 89.9 (0.9)
suSPT 75.4 (1.4) 76.1 (0.8) 92.3 (0.5) 76.4 (1.1) 76.4 (1.2) 88.8 (0.6)
Pooling 82.7 82.9 95.9 74.9 74.2 90.4
HiDisc ssSPT 84.2 (1.4) 83.9 (1.2) 95.6 (0.4) 77.5 (1.3) 77.1 (1.3) 91.2 (0.3)
suSPT 87.4 (0.8) 87.5 (0.7) 93.3 (0.3) 83.6 (0.3) 82.4 (0.2) 88.3 (0.3)
Pooling 55.1 56.4 86.6 54.7 55.0 77.7
HIPT ssSPT 70.1 (0.4) 71.9 (0.5) 92.2 (0.4) 74.9 (1.7) 73.9 (1.7) 89.2 (0.2)
suSPT 72.5 (1.1) 73.7 (0.9) 90.9 (0.3) 76.6 (1.8) 76.1 (1.6) 90.6 (0.7)
Pooling 51.7 51.9 86.8 66.8 66.3 88.0
PLIP ssSPT 68.2 (0.4) 69.7 (0.4) 92.0 (0.4) 76.2 (2.0) 75.5 (2.2) 91.4 (0.4)
suSPT 72.1 (1.4) 72.6 (1.2) 91.3 (0.4) 82.6 (0.7) 81.5 (0.9) 92.5 (0.2)
Pooling 71.1 73.4 93.6 80.3 78.3 93.5
UNI ssSPT 84.8 (1.7) 85.7 (1.5) 96.2 (0.3) 82.8 (2.1) 82.3 (2.2) 94.2 (0.5)
suSPT 86.5 (0.9) 86.9 (0.6) 92.4 (0.5) 89.1 (0.9) 88.8 (0.7) 92.5 (0.5)
Table 9: SPT results with different patch encoder. Metrics reported for results in Figure 5(b). Standard deviations are in (parentheses).
C.4 suSPT benchmarks with UNI patch features
We evaluated suSPT with UNI patch features in Table 3. Error bars are reported in Table 10.
kNN Linear evaluation
Benchmark Method MCA F1 AUC MCA F1 AUC
SRH CNS ABMIL 84.1 (0.9) 84.7 (0.8) 95.4 (0.2) 86.3 (0.9) 86.3 (0.7) 96.3 (0.0)
suSPT 86.5 (0.9) 86.9 (0.6) 92.4 (0.5) 86.4 (0.8) 86.8 (0.5) 98.4 (0.3)
H&E Glioma ABMIL 86.4 (1.5) 85.6 (1.4) 94.7 (1.1) 86.8 (0.6) 85.2 (0.8) 95.5 (0.6)
suSPT 89.1 (0.9) 88.8 (0.7) 92.5 (0.5) 89.1 (1.0) 88.7 (0.7) 97.3 (0.3)
TCGA BRCA ABMIL 85.7 (6.8) 78.8 (9.5) 91.7 (4.7) 88.6 (6.3) 79.1 (6.5) 88.8 (7.6)
suSPT 90.1 (6.4) 84.2 (9.8) 90.3 (5.8) 89.6 (6.3) 82.2 (9.2) 94.7 (3.5)
TCGA NSCLC ABMIL 95.6 (1.2) 95.7 (1.1) 97.3 (2.0) 93.6 (1.7) 90.6 (2.0) 93.4 (2.9)
suSPT 95.4 (2.4) 95.5 (2.3) 95.9 (1.8) 95.5 (2.2) 95.6 (2.2) 97.6 (2.1)
TCGA RCC ABMIL 95.4 (1.3) 93.4 (2.4) 98.3 (0.7) 95.2 (1.1) 93.3 (2.0) 98.5 (0.9)
suSPT 96.3 (2.6) 94.9 (2.9) 97.2 (1.7) 96.2 (2.8) 94.9 (3.1) 98.9 (0.6)
Table 10: suSPT benchmark with UNI features. Extended results for Table 3 with standard deviations reported in (parentheses).
C.5 Ablation stuides
In this section, we present ablation studies over SSL paradigms, SPT transformation choices, and SPT transformation parameters.
C.5.1 SSL paradigms
We examine the effect of different SSL paradigms on ssSPT in table 11. VICReg performs the best on the SRH CNS benchmark, BYOL performs the best on the TCGA RCC benchmark, while SimCLR performs the best on H&E glioma, TCGA BRCA, and TCGA NSCLC. These results demonstrate that the flexibility of SPT allows different algorithms to accommodate different training dynamics of different diagnostic tasks and image modalities.
kNN Linear evaluation
Benchmark Algorithm MCA F1 AUC MCA F1 AUC
ssSPT-SimCLR 77.3 (0.6) 77.6 (0.3) 94.0 (0.4) 82.7 (0.7) 81.8 (0.5) 97.5 (0.3)
SRH CNS ssSPT-VICReg 82.3 (0.5) 82.3 (0.3) 94.7 (0.1) 85.4 (0.3) 83.2 (0.3) 97.6 (0.3)
ssSPT-BYOL 79.4 (0.4) 78.7 (1.2) 94.3 (0.4) 83.4 (0.5) 81.0 (1.3) 96.9 (0.0)
ssSPT-SimCLR 76.5 (0.5) 76.1 (0.4) 90.9 (0.4) 80.0 (2.4) 78.0 (2.2) 91.9 (0.8)
H&E Glioma ssSPT-VICReg 72.8 (0.7) 72.4 (0.9) 87.9 (0.7) 74.5 (0.8) 73.1 (0.8) 90.9 (0.5)
ssSPT-BYOL 74.6 (0.9) 73.8 (0.9) 89.0 (0.5) 77.8 (0.6) 76.1 (0.8) 92.2 (0.0)
ssSPT-SimCLR 72.7 (4.9) 58.3 (6.8) 80.4 (7.8) 83.0 (7.3) 67.4 (9.9) 89.4 (6.3)
TCGA BRCA ssSPT-VICReg 69.2 (7.9) 51.0 (15.2) 77.4 (9.3) 78.6 (5.4) 59.3 (7.5) 85.3 (7.7)
ssSPT-BYOL 70.0 (8.3) 52.6 (15.8) 81.1 (9.0) 83.6 (5.4) 66.7 (6.7) 90.4 (6.6)
ssSPT-SimCLR 86.1 (2.3) 86.1 (2.6) 92.3 (1.7) 88.0 (3.1) 88.0 (3.2) 94.8 (2.1)
TCGA NSCLC ssSPT-VICReg 80.2 (3.4) 80.7 (3.0) 86.0 (3.4) 87.6 (2.2) 87.9 (2.0) 93.8 (1.6)
ssSPT-BYOL 81.4 (2.4) 81.8 (2.7) 87.2 (2.0) 86.0 (2.8) 86.2 (3.1) 92.7 (2.2)
ssSPT-SimCLR 90.7 (2.9) 90.4 (2.7) 98.3 (1.0) 91.7 (2.1) 89.4 (3.2) 98.2 (0.7)
TCGA RCC ssSPT-VICReg 87.2 (3.9) 87.2 (3.1) 96.0 (1.6) 88.6 (3.3) 86.6 (3.2) 96.9 (0.9)
ssSPT-BYOL 91.7 (2.1) 90.5 (2.7) 98.3 (1.3) 93.4 (1.7) 91.1 (2.1) 98.7 (1.1)
Table 11: SPT benchmark with different SSL paradigms. We compare different SPT training algorithms. Standard deviations are reported in (parentheses).
C.5.2 SPT transformation choices
Table 12 reports the performance of individual and combinations of slide-level transformations. We leave out splitting alone since it does not reduce the total number of patch tokens for both views, making it prohibitive to train. For the masking and cropping alone, we show that cropping performs better than masking for both SRH CNS and H&E Glioma. Combining cropping and masking shows similar performance with cropping alone, but is more efficient to train due to the reduced number of tokens per view. Combining each transformation with splitting improves or maintains performance, showing the benefit of further decreasing mutual information (MI) between views. Overall, cropping is the most important transformation, because capturing regional heterogeneity is the most effective and challenging pre-text task.
SRH CNS H&E Glioma
Splitting Cropping Masking MCA F1 MCA F1
X 75.4 (0.9) 76.0 (0.7) 69.9 (0.5) 69.7 (0.5)
X 81.3 (1.5) 81.5 (1.5) 74.3 (0.2) 74.4 (0.2)
X X 76.1 (0.3) 76.3 (0.8) 69.2 (1.5) 69.3 (1.5)
X X 82.1 (1.4) 82.2 (1.2) 73.6 (0.6) 73.8 (0.4)
X X 81.2 (0.5) 81.8 (0.5) 75.5 (2.1) 75.4 (2.2)
X X X 82.3 (0.5) 82.3 (0.3) 76.5 (0.5) 76.1 (0.4)
Table 12: Ablation over transformations. X stands for transformations used. Standard deviations reported in (parentheses).
C.5.3 SPT transformation parameters
We examine the relationship between mutual information (MI) and model performance by combining and adjusting SPT transformation parameters in the H&E glioma benchmark.
Token limit.
We studied the effect of the number of tokens, with splitting + cropping + masking and only masking. A higher token limit results in a higher likelihood of overlapping area coverage between views. The results are in Figure 9a. As expected, MI levels can be adjusted to optimize model performance. Previous studies have concluded that fewer patch tokens resulted in better model performance [15], but only in the case of using masking transformation alone.
Cropping size.